OpenStudy (anonymous):

The equation of line EF is y = 1/2x + 6. Write an equation of a line parallel to line EF in slope-intercept form that contains point (0, −2).

OpenStudy (anonymous):

y = −2x − 2 y = -1/2x + 2 y = 1/2x − 2 y = − 2x + 2

OpenStudy (anonymous):

@jim_thompson5910 @cwrw238 @Abhisar Any help?

jimthompson5910 (jim_thompson5910):

what is the slope of your given line?

jimthompson5910 (jim_thompson5910):

what is the slope of line EF?

OpenStudy (anonymous):

the slope is 1/2, right? @jim_thompson5910

OpenStudy (anonymous):

@jim_thompson5910 you there?

jimthompson5910 (jim_thompson5910):

sorry just got these notifications

jimthompson5910 (jim_thompson5910):

yes the slope of EF is 1/2

jimthompson5910 (jim_thompson5910):

ANY other line will also have a slope of 1/2 why? because any two parallel lines have equal slopes (just different y intercepts)

jimthompson5910 (jim_thompson5910):

so ANY line parallel to EF is of this form y = (1/2)x + b where b is some fixed number

OpenStudy (anonymous):

okaay im starting to understand..

jimthompson5910 (jim_thompson5910):

let's force the line to go through (0,-2) y = (1/2)x + b -2 = (1/2)(0) + b -2 = 0 + b -2 = b b = -2 we know that m = 1/2 and b = -2 so the equation is y = (1/2)x - 2

jimthompson5910 (jim_thompson5910):

notice how I plugged in (x,y) = (0,-2)

OpenStudy (anonymous):

Ohhhh! you are a life saver! Thank you

OpenStudy (anonymous):

@jim_thompson5910 can you help with one more? I'll medal after.

jimthompson5910 (jim_thompson5910):

sure

OpenStudy (anonymous):

The equation of line AB is (y−3) = 5 (x − 4). What is the slope of a line perpendicular to line AB? −5 5 -1/5 1/5

jimthompson5910 (jim_thompson5910):

what is the slope of line AB? You'll have to solve for y to determine this.

jimthompson5910 (jim_thompson5910):

or nvm, you can compare that to y - y1 = m(x-x1) to determine the slope m

OpenStudy (anonymous):

I'm not understanding..

jimthompson5910 (jim_thompson5910):